# Log Probability Tracking of LLM APIs

### v1

@@ 1 Introduction
LLM API providers typically offer version-pinned endpoints, signaling to users that a given endpoint will serve a consistent model. Users of APIs tend to rely on this consistency: developers want to avoid unexpected regressions in their applications; researchers seek reproducibility in their experiments; regulators perform initial compliance assessments, and assume that the API will keep serving the same model afterward (yan2022active). Yet users have no practical way to verify this consistency. Once a niche concern, this gap is drawing more attention as LLMs get deployed in mission-critical applications, notably in software engineering.
LLMs are, however, highly complex software artifacts under constant development. LLM API providers may change their inference software and hardware infrastructure for performance reasons, or modify the model itself in response to new jailbreaks or to update its behavior. Some providers might also quietly deploy quantized versions to save costs, switch to more light-weight models during peak traffic, or even serve different variants based on query characteristics. More concerning from a security perspective, providers might suffer the malicious injection of unobstrusive model backdoors, infringing the privacy and security of their users for an attacker’s benefit. Such explicit attacks could happen from insiders, a bad update, or a compromised supply chain. This concern isn’t just theoretical: the version of Grok deployed on X had three such incidents in 2025 where a modified system prompt was deployed to users of the X platform. Two incidents were blamed on rogue employees (grokfeb; grokmay) and one on a bad update (grokjuly).
Unfortunately, existing change-detection methods tend to be prohibitively expensive. State of the art approaches typically rely on the processing of tokens returned by the monitored LLM (chen2023chatgptsbehaviorchangingtime; DBLP:conf/iclr/GaoLG25; cai2025gettingpayforauditing) and require extensive benchmarking or statistical analysis across many queries to reach a robust conclusion. As a result, LLM APIs are left largely unmonitored by third parties for changes in practice.
@@
Intuitively, logprobs constitute a far richer information source than tokens, yielding a handful of continuous values along with a token. Unfortunately, exploiting these logprobs is not trivial. Logprobs are not deterministic in practice (cai2025gettingpayforauditing), making them more challenging to use for model change auditing than simply checking logprob vectors for equality across time.
@@
All our code is open-source at https://github.com/timothee-chauvin/track-llm-apis.
@@ 2.1 Rationale
We follow the definition of the model equality testing problem in DBLP:conf/iclr/GaoLG25: we consider a pair of LLM APIs (which can be the same API at different times).
@@
These logprobs encode a part of the probability distribution, and as such are denser in information than the sampled tokens. For instance, they can be used to recover the prompt tokens given an output (morris2023languagemodelinversion), and have been used to leak proprietary information about an LLM, including its hidden dimension and embedding projection layer (finlayson2024logitsapiprotectedllmsleak; carlini2024stealingproductionlanguagemodel).
This motivates us to study the feasibility of detecting changes in LLMs via the top-$k$ logprobs of a single output token, on a fixed short prompt. Intuitively, detecting many types of change using only a single prompt and a single output token seems possible, as LLM modifications, such as finetuning or quantization, typically affect all the weights of a model. In particular, finetuning in one domain is known to affect completions in other domains (qi2024finetuning).
@@ 2.2 Non-determinism in LLMs
  • Unintentional non-determinism: In practice however, logits often fluctuate between identical requests, resulting in a non-deterministic logprob vector. While the majority of inference kernels are run-to-run deterministic at the batch level, individual requests are non-deterministic, as they are affected by other requests in the batch (he2025nondeterminism). In addition, the software and hardware stack on top of which the model is deployed is also a source of non-determinism in production LLM APIs, where successive requests are likely to be routed to different GPUs (zhang2024hardwaresoftwareplatforminference).
@@ 3.1 Evaluation Setup
This benchmark takes an LLM as input and generates a set of 58 variants, ranging from minor to more substantial modifications, reflecting practical operations a model might sustain in practice. Variants are produced across five levels of modification intensity, with difficulty increasing in powers of two. Specifically, we apply: regular fine-tuning and LoRA fine-tuning [^1] A parameter-efficient fine-tuning method that introduces trainable low-rank matrices into pre-trained weights, enabling effective adaptation with far fewer trainable parameters (hu2022lora), each for 1 epoch with between 1 and 512 single-sample finetuning steps; unstructured weight pruning, either by magnitude or by random selection, removing a fraction of weights ranging from $2^{-10}$ up to 1; and parameter noising, adding independent Gaussian noise to each parameter with standard deviation $\sigma$ ranging from $2^{-15}$ to 1. More details in Appendix A.2.
The fine-tuning data is sampled from the LMSYS-Chat-1M dataset (zheng2024lmsyschatm), a dataset of real-world conversations between users and multiple LLMs, selecting only single-turn interactions with GPT-4. The goal is for the finetuning dataset to be both difficult to detect (because the data is in-distribution for models similar to GPT-4) and realistic (because most changes to production models are likely to be subtle).
We apply the TinyChange benchmark to 5 open-weight LLMs from 0.5B to 8B parameters: Qwen 2.5 0.5B, Gemma 3 1B, Phi-3 Mini 4k, Llama 3.1 8B, and OLMo 2 7B [^2] Huggingface IDs: Qwen/Qwen2.5-0.5B-Instruct, google/gemma-3-1b-it, microsoft/Phi-3-mini-4k-instruct, meta-llama/Llama-3.1-8B-Instruct, and allenai/OLMo-2-1124-7B-Instruct, resulting in a total of 290 variants.
@@
  • MET (Model Equality Testing method, DBLP:conf/iclr/GaoLG25) uses Maximum Mean Discrepancy (MMD) with a Hamming distance kernel to compare two distributions of responses. As in the original paper, we use $P_{MET}=25$ prompts, 50-token responses, and $T=1$.
  • MMLU-ALG MMLU (Massive Multitask Language Understanding, hendryckstest2021) is a well-known benchmark of the knowledge of LLMs. As the benchmark contains 14,042 questions, we restrict it to the “abstract_algebra” subset with $P_{MA}=100$ questions, and call this baseline MMLU-ALG. Using MMLU is similar to the approach used in chen2023chatgptsbehaviorchangingtime, with the difference that MMLU consists of questions with multiple-choice answers A/B/C/D, thus considerably driving down the token cost of this approach compared to freeform answers. Each of the 100 prompts is sampled at $T=0.1$, generating 5 output tokens. Similarly to 2.3, we define the two-sample test statistic as the mean absolute distance between per-question accuracies. More details in Appendix A.3.
@@
For each of the 5 original models and 290 variants ($V=295$), we generate $S=10,000$ samples per prompt using vLLM (kwon2023efficient). The MET and MMLU-ALG prompts are sampled at $T>0$, where the non-determinism from temperature sampling far outweighs unintentional non-determinism. For LT however, we are concerned with the logprobs, so we introduce realistic noise from other requests by mixing our prompts with random traffic, also sampled from LMSYS-Chat-1M for simplicity. To generate each of the 10,000 samples, we mix and shuffle our $P_{LT}=16$ tested prompts with 48 random prompts, and generate the logprobs for the first token of the batch with vLLM. (LT uses a single prompt, but we evaluated the influence of the prompt in 3.2.1). This results in $V\times S\times(P_{MET}+P_{MA}+P_{LT})=416\text{M}$ outputs.
@@ 4 Limitations and future work
A provider may also refuse to serve requests with only one or a few tokens of output. Indeed, we found that between 16:00 and 17:00 UTC on September 24, 2025, OpenAI started requiring at least 16 tokens of output on their GPT-4.1 line of models accessed via OpenRouter (though not on their own API)333at least for our one-token requests that requested output logprobs. The error message is: ”Invalid ’max_output_tokens’: integer below minimum value. Expected a value $\geq$ 16, but got 1 instead.”.
@@
We showed that using logprobs with a simple hypothesis testing method enables extremely sensitive change detection ”out of the box”, but detection performance can likely be increased further by tuning parameters. While the hypothesis testing framework simple, applicable, and enables comparison with other methods, alternative approaches such as CUSUM (cusum) control charts could also be explored.
@@ 5 Related Work
Prior to widespread LLM deployment, chen2022how introduced methods to assess ML API shifts, and found significant shifts in a third of the ML APIs they tested.
@@
chen2023chatgptsbehaviorchangingtime compared versions of ChatGPT using multiple benchmarks, but this approach requires extensive testing across thousands of queries, and the versions they tested were explicitly not the same version. DailyBench (phillips2025dailybench) was an actual attempt at benchmark-based continuous monitoring of LLM APIs, testing 5 LLMs 4 times a day on a version of HELMLite (liang2023holistic), but it was discontinued after 40 days due to costing \$5 a day. DBLP:conf/iclr/GaoLG25 compared output distributions with the Maximum Mean Discrepancy statistical test, using a Hamming distance kernel. cai2025gettingpayforauditing re-implemented and compared multiple detection techniques.
@@
Change detection and fingerprinting are closely related, as it’s often possible to convert a change detection method into a fingerprinting method, and vice versa. However, some fingerprinting methods explicitly aim to be invariant to small changes in a model, or only attempt to identify an LLM within a closed set of known models. pasquini2025llmmapfingerprintinglargelanguage introduced a set of handwritten queries coupled with a classifier to identify which known model a given LLM is closest to. They explicitly aimed to produce queries that would be insensitive to small changes in a model, but sensitive to large changes. sun2025idiosyncrasieslargelanguagemodels fine-tuned embedding models on LLM-generated texts, to identify which LLM out of a set of 5 generated a given text.
@@
he2025nondeterminism demonstrated that run-to-run (i.e. on the same GPU) LLM non-determinism is due to a changing batch size based on the load on the inference server, proposing batch-invariant kernels as a solution. zhang2024hardwaresoftwareplatforminference leveraged within-platform determinism and cross-platform non-determinism to identify which hardware and software configuration a model is running on, also proposing a method based on logprobs.
@@
Formal verification is another approach for the problem we are concerned with: if an inference result can be proven to come from a specific model, then we can be confident that the model has not changed. zkLLM (10.1145/3658644.3670334) introduced zero-knowledge proofs for LLMs. However, the proofs take several minutes to generate for a 13b LLM, and are on the order of 200kB in size, making this approach prohibitively expensive for high-performance inference servers. TOPLOC (ong2025toploclocalitysensitivehashing) used locality-sensitive hashing to considerably reduce the overhead of proving an inference result, but provider cooperation is still required.

### v2

# arXiv:2512.03816v2[cs.LG] 27 Feb 2026

@@ 1 Introduction
footnotetext: Code available at: https://github.com/timothee-chauvin/track-llm-apis
LLM API providers typically offer version-pinned endpoints, signaling to users that a given endpoint will serve a consistent model. Users of APIs tend to rely on this consistency: developers want to avoid unexpected regressions in their applications; researchers seek reproducibility in their experiments; regulators perform initial compliance assessments, and assume that the API will keep serving the same model afterward (Yan and Zhang, 2022). Yet users have no practical way to verify this consistency. Once a niche concern, this gap is drawing more attention as LLMs get deployed in mission-critical applications, notably in software engineering.
LLMs are, however, highly complex software artifacts under constant development. LLM API providers may change their inference software and hardware infrastructure for performance reasons, or modify the model itself in response to new jailbreaks or to update its behavior. Some providers might also quietly deploy quantized versions to save costs, switch to more light-weight models during peak traffic, or even serve different variants based on query characteristics. More concerning from a security perspective, providers might suffer the malicious injection of unobstrusive model backdoors, infringing the privacy and security of their users for an attacker’s benefit. Such explicit attacks could happen from insiders, a bad update, or a compromised supply chain. This concern isn’t just theoretical: the version of Grok deployed on X had three such incidents in 2025 where a modified system prompt was deployed to users of the X platform. Two incidents were blamed on rogue employees (Babuschkin, ; xAI, ) and one on a bad update (xAI, ).
Unfortunately, existing change-detection methods tend to be prohibitively expensive. State of the art approaches typically rely on the processing of tokens returned by the monitored LLM (Chen et al., 2023; Gao et al., 2025; Cai et al., 2025) and require extensive benchmarking or statistical analysis across many queries to reach a robust conclusion. As a result, LLM APIs are left largely unmonitored by third parties for changes in practice.
@@
Intuitively, logprobs constitute a far richer information source than tokens, yielding a handful of continuous values along with a token. Unfortunately, exploiting these logprobs is not trivial. Logprobs are not deterministic in practice (Cai et al., 2025), making them more challenging to use for model change auditing than simply checking logprob vectors for equality across time.
@@ 2.1 Rationale
We follow the definition of the model equality testing problem in Gao et al. (2025): we consider a pair of LLM APIs (which can be the same API at different times).
@@
These logprobs encode a part of the probability distribution, and as such are denser in information than the sampled tokens. For instance, they can be used to recover the prompt tokens given an output (Morris et al., 2023), and have been used to leak proprietary information about an LLM, including its hidden dimension and embedding projection layer (Finlayson et al., 2024; Carlini et al., 2024).
This motivates us to study the feasibility of detecting changes in LLMs via the top-$k$ logprobs of a single output token, on a fixed short prompt. Intuitively, detecting many types of change using only a single prompt and a single output token seems possible, as LLM modifications, such as finetuning or quantization, typically affect all the weights of a model. In particular, finetuning in one domain is known to affect completions in other domains (Qi et al., 2024).
@@ 2.2 Non-determinism in LLMs
  • Unintentional non-determinism: In practice however, logits often fluctuate between identical requests, resulting in a non-deterministic logprob vector. While the majority of inference kernels are run-to-run deterministic at the batch level, individual requests are non-deterministic, as they are affected by other requests in the batch (He and Lab, 2025). In addition, the software and hardware stack on top of which the model is deployed is also a source of non-determinism in production LLM APIs, where successive requests are likely to be routed to different GPUs (Zhang et al., 2024).
@@ 3.1 Evaluation Setup
This benchmark takes an LLM as input and generates a set of 58 variants, ranging from minor to more substantial modifications, reflecting practical operations a model might sustain in practice. Variants are produced across five levels of modification intensity, with difficulty increasing in powers of two. Specifically, we apply: regular fine-tuning and LoRA fine-tuningA parameter-efficient fine-tuning method that introduces trainable low-rank matrices into pre-trained weights, enabling effective adaptation with far fewer trainable parameters (Hu et al., 2022), each for 1 epoch with between 1 and 512 single-sample finetuning steps; unstructured weight pruning, either by magnitude or by random selection, removing a fraction of weights ranging from $2^{-10}$ up to 1; and parameter noising, adding independent Gaussian noise to each parameter with standard deviation $\sigma$ ranging from $2^{-15}$ to 1. More details in Appendix A.2.
The fine-tuning data is sampled from the LMSYS-Chat-1M dataset (Zheng et al., 2024), a dataset of real-world conversations between users and multiple LLMs, selecting only single-turn interactions with GPT-4. The goal is for the finetuning dataset to be both difficult to detect (because the data is in-distribution for models similar to GPT-4) and realistic (because most changes to production models are likely to be subtle).
We apply the TinyChange benchmark to 5 open-weight LLMs from 0.5B to 8B parameters: Qwen 2.5 0.5B, Gemma 3 1B, Phi-3 Mini 4k, Llama 3.1 8B, and OLMo 2 7BHuggingface IDs: Qwen/Qwen2.5-0.5B-Instruct, google/gemma-3-1b-it, microsoft/Phi-3-mini-4k-instruct, meta-llama/Llama-3.1-8B-Instruct, and allenai/OLMo-2-1124-7B-Instruct, resulting in a total of 290 variants.
@@
  • MET (Model Equality Testing method, Gao et al. (2025)) uses Maximum Mean Discrepancy (MMD) with a Hamming distance kernel to compare two distributions of responses. As in the original paper, we use $P_{MET}=25$ prompts, 50-token responses, and $T=1$.
  • MMLU-ALG MMLU (Massive Multitask Language Understanding, Hendrycks et al. (2021)) is a well-known benchmark of the knowledge of LLMs. As the benchmark contains 14,042 questions, we restrict it to the “abstract_algebra” subset with $P_{MA}=100$ questions, and call this baseline MMLU-ALG. Using MMLU is similar to the approach used in Chen et al. (2023), with the difference that MMLU consists of questions with multiple-choice answers A/B/C/D, thus considerably driving down the token cost of this approach compared to freeform answers. Each of the 100 prompts is sampled at $T=0.1$, generating 5 output tokens. Similarly to 2.3, we define the two-sample test statistic as the mean absolute distance between per-question accuracies. More details in Appendix A.3.
@@
For each of the 5 original models and 290 variants ($V=295$), we generate $S=10,000$ samples per prompt using vLLM (Kwon et al., 2023). The MET and MMLU-ALG prompts are sampled at $T>0$, where the non-determinism from temperature sampling far outweighs unintentional non-determinism. For LT however, we are concerned with the logprobs, so we introduce realistic noise from other requests by mixing our prompts with random traffic, also sampled from LMSYS-Chat-1M for simplicity. To generate each of the 10,000 samples, we mix and shuffle our $P_{LT}=16$ tested prompts with 48 random prompts, and generate the logprobs for the first token of the batch with vLLM. (LT uses a single prompt, but we evaluated the influence of the prompt in 3.2.1). This results in $V\times S\times(P_{MET}+P_{MA}+P_{LT})=416\text{M}$ outputs.
@@ 4 Limitations and future work
A provider may also refuse to serve requests with only one or a few tokens of output. Indeed, we found that between 16:00 and 17:00 UTC on September 24, 2025, OpenAI started requiring at least 16 tokens of output on their GPT-4.1 line of models accessed via OpenRouter (though not on their own API)at least for our one-token requests that requested output logprobs. The error message is: ”Invalid ’max_output_tokens’: integer below minimum value. Expected a value $\geq$ 16, but got 1 instead.”.
@@
We showed that using logprobs with a simple hypothesis testing method enables extremely sensitive change detection ”out of the box”, but detection performance can likely be increased further by tuning parameters. While the hypothesis testing framework simple, applicable, and enables comparison with other methods, alternative approaches such as CUSUM (Page, 1954) control charts could also be explored.
@@ 5 Related Work
Prior to widespread LLM deployment, Chen et al. (2022) introduced methods to assess ML API shifts, and found significant shifts in a third of the ML APIs they tested.
@@
Chen et al. (2023) compared versions of ChatGPT using multiple benchmarks, but this approach requires extensive testing across thousands of queries, and the versions they tested were explicitly not the same version. DailyBench (Phillips, 2025) was an actual attempt at benchmark-based continuous monitoring of LLM APIs, testing 5 LLMs 4 times a day on a version of HELMLite (Liang et al., 2023), but it was discontinued after 40 days due to costing \$5 a day. Gao et al. (2025) compared output distributions with the Maximum Mean Discrepancy statistical test, using a Hamming distance kernel. Cai et al. (2025) re-implemented and compared multiple detection techniques.
@@
Change detection and fingerprinting are closely related, as it’s often possible to convert a change detection method into a fingerprinting method, and vice versa. However, some fingerprinting methods explicitly aim to be invariant to small changes in a model, or only attempt to identify an LLM within a closed set of known models. Pasquini et al. (2025) introduced a set of handwritten queries coupled with a classifier to identify which known model a given LLM is closest to. They explicitly aimed to produce queries that would be insensitive to small changes in a model, but sensitive to large changes. Sun et al. (2025) fine-tuned embedding models on LLM-generated texts, to identify which LLM out of a set of 5 generated a given text.
@@
He and Lab (2025) demonstrated that run-to-run (i.e. on the same GPU) LLM non-determinism is due to a changing batch size based on the load on the inference server, proposing batch-invariant kernels as a solution. Zhang et al. (2024) leveraged within-platform determinism and cross-platform non-determinism to identify which hardware and software configuration a model is running on, also proposing a method based on logprobs.
@@
Formal verification is another approach for the problem we are concerned with: if an inference result can be proven to come from a specific model, then we can be confident that the model has not changed. zkLLM (Sun et al., 2024) introduced zero-knowledge proofs for LLMs. However, the proofs take several minutes to generate for a 13b LLM, and are on the order of 200kB in size, making this approach prohibitively expensive for high-performance inference servers. TOPLOC (Ong et al., 2025) used locality-sensitive hashing to considerably reduce the overhead of proving an inference result, but provider cooperation is still required.
@@ 7 Reproducibility statement
8 Acknowledgments
We acknowledge the support of the French Agence Nationale de la Recherche (ANR), under grant ANR-24-CE23-7787 (project PACMAM).
This work was supported by the Cluster SequoIA Chair FANG funded by ANR, reference number ANR-23-IACL-0009.
This project was provided with computing (AI) and storage resources by GENCI at IDRIS thanks to the grant 2025-AD011016369 on the supercomputer Jean Zay’s H100 partition.